# FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with Multimodal Large Language Models

### v1

@@
FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with Multimodal Large Language Models
Hongyang Wang1, Yichen Shi2, Zhuofu Tao3, Yuhao Gao1, Liepiao Zhang4, Xun Lin5
Jun Feng1, Xiaochen Yuan6, Zitong Yu5🖂, Xiaochun Cao7
1Shijiazhuang Tiedao University   2Shanghai Jiao Tong University  3UCLA  4GRGBanking
5Great Bay University   6Macao Polytechnic University   7Shenzhen Campus of Sun Yat-sen University
Equal contribution      🖂 Corresponding author
Figure 1: FaceShield Multi-task Response Demonstration. This figure illustrates the model’s performance across four tasks: coarse-grained classification, fine-grained classification, reasoning, and localization. Coarse-grained classification distinguishes real from spoofed faces, while fine-grained classification identifies specific attack types, such as print attacks. The reasoning task provides explanations for spoofing by analyzing features like lighting and symmetry. Localization precisely detects spoofed regions. This illustration highlights FaceShield’s capability to handle a range of complex questions with accuracy.
Face anti-spoofing (FAS) is crucial for protecting facial recognition systems from presentation attacks. Previous methods approached this task as a classification problem, lacking interpretability and reasoning behind the predicted results. Recently, multimodal large language models (MLLMs) have shown strong capabilities in perception, reasoning, and decision-making in visual tasks. However, there is currently no universal and comprehensive MLLM and dataset specifically designed for FAS task. To address this gap, we propose FaceShield, a MLLM for FAS, along with the corresponding pre-training and supervised fine-tuning (SFT) datasets, FaceShield-pre10K and FaceShield-sft45K. FaceShield is capable of determining the authenticity of faces, identifying types of spoofing attacks, providing reasoning for its judgments, and detecting attack areas. Specifically, we employ spoof-aware vision perception (SAVP) that incorporates both the original image and auxiliary information based on prior knowledge. We then use an prompt-guided vision token masking (PVTM) strategy to random mask vision tokens, thereby improving the model’s generalization ability. We conducted extensive experiments on three benchmark datasets, demonstrating that FaceShield significantly outperforms previous deep learning models and general MLLMs on four FAS tasks, i.e., coarse-grained classification, fine-grained classification, reasoning, and attack localization. Our instruction datasets, protocols, and codes will be released soon.
1 Introduction
@@
Existing deep learning FAS models can be categorized into two types: vision-based methods and vision-language-based methods. As shown in Fig. 2(a), vision-based methods rely solely on image data (e.g., RGB, Depth, Infrared(IR)) and binary labels to train CNNs [66, 67] or ViTs [15, 70] for FAS. While they can achieve satisfactory results against known attack types and environments, these methods are prone to overfitting on spurious correlations and lack strong extrapolation capabilities. As illustrated in Fig. 2(b), Vision-language-based methods do not use binary labels but instead train CLIPs with image-text pairs [55, 36, 43]. The text labels in these methods provide more domain-agnostic information, enhancing models’ generalization capability. Although these FAS models demonstrate some recognition capabilities, they still face challenges such as limited generalization ability, poor interpretability, and a lack of capability for fine-grained localization of attack regions.
Recently, MLLMs have shown remarkable capabilities across various visual tasks, such as remote sensing [76, 29, 44], medical imaging [32, 57], and deepfake detection  [64, 25]. By leveraging the general capabilities of language foundation models alongside the visual information extracted by vision towers, these specialized MLLMs integrate perception, reasoning, and decision-making within a single model. Regarding the FAS task, SHIELD [53] conducted extensive evaluations on existing general-purpose MLLMs, revealing that their performance on FAS tasks still has room for improvement. [75] introduced a model capable of performing classification and description attack type. However, the model is limited in its ability to handle more nuanced tasks, such as identifying specific attack types, reasoning, and localizing spoofed areas. These limitations underscore the broader challenges in training FAS MLLMs, including: (1) a lack of pretraining and SFT datasets specific to FAS tasks, (2) the need to extend traditional FAS tasks to fully exploit MLLM capabilities, and (3) the difficulty for general-purpose vision towers to capture the subtle distinctions between real faces and PAs, unlike with natural images.
@@ 2 Related Work
Figure 3: Construction pipeline of our proposed instruction datasets (i.e., FaceShield-pre10K and FaceShield-sft45K).The initial datasets (WMCA, PADISI, SiW-Mv2) are combined to form a uni-class dataset covering 12 spoofing types, with selected images annotated for visual grounding. Using MLLM with structured prompts, we generate two datasets: a pretraining dataset and an SFT dataset divided into four tasks (coarse-grained classification, fine-grained classification, reasoning, and localization). The pretraining data is filtered by CLIP for similarity, producing the FaceShield-pre10k dataset. SFT data undergoes multi-level filtering (LLM-based, keywords, and human reviews), followed by augmentation, resulting in the FaceShield-sft45k dataset.
Face Anti-Spoofing. Previous works usually employed convolutional neural networks (CNNs) [66, 67] or vision transformers (ViTs) [15, 70, 10] as backbone networks, integrating auxiliary priors for live/spoof classification (e.g., reflection light [6], depth reconstruction [6, 41, 3], rPPG signals [33, 65, 41, 73], and gradient information [68, 8]) to detect PAs. While these methods perform well in source domains, their performance degrades significantly in unseen domains due to domain shifts and unseen attacks [71]. Recently, many domain-generalized FAS methods [71] have been proposed to address these cross-domain challenges. Adversarial training [27, 26] and feature disentanglement techniques [42, 28, 17, 62] have been explored to decouple image features into domain-invariant (content) and domain-specific (style) features, preventing the model from learning spurious correlations between domain-specific features and labels. Some approaches employ meta-learning [48, 52, 7], data augmentation [9], or data synthesis methods [14, 39] to simulate domain shifts, enabling the model to learn a broader data distribution during training. Others perform domain [26, 58, 21, 60, 78] or gradient [30] alignment to constrain the relative relationships between sample embeddings across different source domains, aiming to learn more generalized decision boundaries. More recently, some methods have adopted vision-language models (VLMs), particularly CLIP [49], to detect PAs using text guidance. By leveraging the reasoning capabilities acquired during large-scale pretraining, VLM-based approaches [56, 36, 61, 37, 40] have shown to improve the generalization of FAS. Unlike traditional binary live/spoof classification, VLM-based methods classify images by computing similarities between visual features and textual descriptions of live and spoof faces, leveraging more causal information [36].
Existing FAS methods lack interpretable reasoning capabilities, e.g., language-based explanations, to distinguish between live and spoof faces. A concurrent work [75] collect 12 datasets and described images based on attack types and keywords. Using instruction tuning, they develop a MLLM capable of classifying and providing detailed descriptions of attacks. Meanwhile, these approaches fail to integrate classification, localization, and reasoning tasks simultaneously, limiting their adaptability and reliability in real-world applications, where versatile understanding is critical for reliable decision-making. Additionally, existing FAS datasets [61, 37] lack annotations for reasoning processes, further hindering the development of reasoning-capable FAS methods.
Figure 4: Comparison of different model architectures. (a) Single modal baseline model. (b) Proposed model without Prompt-Guided Vision Token Masking (PVTM). (c) Proposed model with PVTM. The input image is first encoded by the vision encoder, and PVTM partitions the encoded features into RGB, SAV, and Masked tokens. These visual tokens are passed to the projector, while the prompt is tokenized separately, producing text tokens. The combined visual and text tokens are then fed into the Large Language Model (LLM) with LoRA, producing the final output.
Multimodal Large Language Model. MLLMs, such as GPT-4V [2], LLaVA [38], Bunny [20] have demonstrated successful performances in general domains [13, 31, 24], attracting considerable attention from researchers. Researchers have developed task-specific MLLMs in various fields under vision tasks, including remote sensing [76, 29, 44], medical [32, 57], aesthetic evaluation [23], deepfake detection [64, 25] and Visual Grounding [4, 63, 46]. Unlike traditional models, MLLMs can handle multiple tasks within a single model and exhibit strong performance in unknown domains. FakeShield [64] proposes an explainable multi-task model for image forgery detection and localization, capable of making judgments based on tampering clues at both the image and pixel levels. The language foundation model contains scarce knowledge of face attacks, resulting in poor performance of existing MLLMs on FAS tasks. Our constructed datasets can introduce abundant FAS-related knowledge to MLLM, which through SAVP further enhances its ability to distinguish between real faces and PAs, and PVTM enhances the generalizability of the extracted features.
3 FaceShield-pre10K and FaceShield-sft45K
3.1 Dataset Collection
@@
Fig. 3(a) shows the collection and localization coordinate annotation process of existing FAS datasets. Based on class types of WMCA (W) [16], PADSIS (P) [51] and SiW-Mv2 (S) [18] datasets, we organize all annotation categories into 12 distinct types: Bonafide, Fakehead, Print, Glasses, Replay, Paper mask, Flexible mask, Rigid mask, Partial Eye, Partial Mouth, Makeup, and Tattoo. We have also re-annotated the original dataset images at both the image and region levels. Finally, we obtained 12,091 images with class labels and 3,139 images with bounding box labels. More details and examples can be found in our Appendix.
3.2 Instruction Construction
After collecting images, class labels, and attack regions, as shown in Fig. 3(b), we use Bunny-Llama-3-8B-V [20] to assist in generating the pretrain dataset and the SFT dataset. Specifically, we first use a system prompt as a preamble to enhance the performance of responses from the MLLM assistant. Class labels are embedded in the prompt, providing descriptions of some image content. Subsequently, given the task type, the MLLM assistant generates multiple question-and-answer pairs for the task, following the format provided in the few-shot examples.
@@ 3.2 Instruction Construction
For the construction of the pretraining dataset FaceShield-pre10K, we only use the MLLM assistant to generate descriptions of the images, without any other information. The image and its corresponding generated description are filtered using a CLIP [50] score, removing image-text pairs with a similarity of less than 15%.
For the task-specific instruction dataset FaceShield-sft45K, after generating question-answer (QA) pairs using MLLM, manual filtering, and specific keyword filtering are applied to remove incorrect QA pairs. We enhance the diversity of the content in questions and answers using Llama3 [12] on the golden seed dataset obtained after filtering, to improve the dialogue capabilities of the model trained on this dataset. The specific content and responses for each task are as follows:
  • Coarse-grained classification: MLLM directly responds with labels for face images, identifying them as either real faces or PAs.
  • Fine-grained classification: The responses from the MLLM include not only real or PAs but also judgments on the specific categories of the PAs.
  • Reasoning: MLLM provides descriptions and reasoning from multiple perspectives on the image and delivers the final judgment.
  • Attack localization: MLLM first evaluates the faces in the image, if PA is detected, it provides the specific coordinates of the attack area.
4 FaceShield
@@
Our goal is to train a FAS task-specific MLLM with two main objectives: 1) Enhance the visual encoder’s ability to extract features from real faces and presentation attacks, and 2) Utilize the extensive knowledge stored in the LLM to improve the model’s generalization capabilities when facing unknown domains. A naive training approach is shown in Fig.  4(a), where direct pre-training and SFT are conducted using RGB images and constructed QA data. However, the high similarity between real faces and PAs in RGB appearance poses significant challenges to this method. As shown in Fig. 4(b), Spoof-Aware Vision Perception (SAVP) combines images preprocessed based on prior knowledge, by extracting predefined local descriptor operators [72], with the original RGB image. Our complete model framework is shown in Fig. 4(c), RGB images and the extracted local descriptor images are fed into the vision encoder to extract vision token $V_{RGB}$ and $V_{SAV}$, respectively. These features are then processed through the Prompt-Guided Vision Token Masking (PVTM) module, which extracts highly generalizable vision tokens. These tokens are sent to a projector to align with text prompt token $P$ to produce $V_{align}$, which is then fed into the language foundation model for inference result $\mathcal{Y}$ as follows:
@@
4.1 Spoof-Aware Vision Perception
As shown in Fig. 4(a), bonafide faces and PAs lack distinct discriminative features in RGB-based appearance space, whereas local descriptors [66, 72] extracted through image preprocessing can enhance their subtle live/spoof clues. As shown in Fig. 4(b), we extract features from the original images using Local Binary Pattern (LBP) [47], Gray, and Histogram of Oriented Gradients (HOG) [11], and concatenate them. LBP and Gray-specific computations are as follows:
@@ 4.1 Spoof-Aware Vision Perception
$s(x)$ =$=\begin{cases}1&\text{if }x\geq 0,\\ 0&\text{otherwise},\end{cases}$
@@
Figure 5: Prompt-guided Vision Token Masking Module (PVTM)
4.2 Prompt-Guided Vision Token Masking
To further enhance the alignment between visual features and text prompts, and alleviate overfitting on spurious correlations, we leverage text prompts to guide token selection after the visual encoder. As shown in Fig. 5, the text tokens extracted from the prompt are pooled [54], and then their similarity with all visual tokens is calculated. We assume that visual tokens with higher similarity are more relevant to subsequent tasks and even for the same image, the important visual tokens may not be consistent across different tasks. The calculation of similarity between visual tokens $V_{i}$ and text prompt tokens $P$ is as follows:
@@ 4.2 Prompt-Guided Vision Token Masking
4.3 Training Details
We employ a two-stage training strategy. The first stage is the pre-training stage, where we perform continual pretrain of existing weights. In this stage, we align the visual embeddings from a pre-trained vision encoder with the text embeddings using FaceShield-pre10K. The second stage is the supervised fine-tuning (SFT) stage, where we apply visual instruction tuning to fully leverage the MLLM’s capabilities across various multimodal tasks within our specific domain using FaceShield-sft45K. For LLM tuning, we incorporate LoRA [22] and set the parameters with a hidden dimension of 128 and a scaling factor of 256 in every transformer block. In both stages, we employ cross-entropy loss for next-token prediction. During the pre-training stage, only the projector after the vision encoder and PVTM are optimized for one epoch. In the SFT stage, we fine-tune the LoRA layers in LLM as well as the projector.
@@ 4.3 Training Details
ResNet [19] 97.55 2.32
PatchNet [59] 98.22 1.78
CoOp [77] 98.73 1.27
@@
LLaVA [38] 65.54 27.76
Qwen-VL [4] 51.94 38.70
Minigpt4 [79] 26.86 65.50
Bunny [20] 81.20 17.87
Bunny (fine-tuned) [20] 98.23 1.52
@@
5 Experiments and Results
5.1 Protocols and Evaluation Metrics
We select 10% of each source dataset in FaceShield-sft45K to form three test subsets (W, S, P). For the coarse-grained classification and fine-grained classificationtask, we conduct both intra-dataset and cross-dataset testing. For the reasoning, and attack localization tasks, we employ intra-dataset evaluation. The intra-dataset test uses all data from the source datasets for training and evaluation on the combined test sets (W&S&P). For cross-dataset testing, we select two datasets for pre-training and SFT, and then test using the test set from the remaining dataset. For example, we might train on W and S, and test using the test set from P. For the coarse-grained classification, fine-grained classification, and reasoning tasks, we use Half Total Error Rate (HTER) [69] or accuracy (ACC) as evaluation metrics. Additionally, we evaluate the reasoning process for the reasoning task using BLEU [45], ROUGE-L [34], and METEOR [5], and evaluate the attack localizationtask using AP@40 and AP@50.
5.2 Implementation Details
We use Siglip [74] as the visual encoder and Phi-3 [1] as the language foundation model. PVTM retains the top 10% of the most important tokens and randomly masks 5% of the tokens in the remaining 90%. Adam optimizer is used in the pretrain stage with a learning rate of $5\times 10^{-4}$. As for SFT stage, we decrease the learning rate to $2\times 10^{-4}$. All experiments are conducted on a single NVIDIA A100 GPU. Each experiment is repeated 10 times on the model, and the final results are reported as the mean ± standard deviation.
@@ 5.2 Implementation Details
W & S P
ResNet [19] 46.12 50.00
PatchNet [59] 77.18 22.87
IADG [78] 72.96 27.01
@@
W & P S
ResNet [19] 53.36 49.16
PatchNet [59] 56.16 45.37
IADG [78] 57.20 42.81
@@
S & P W
ResNet [19] 74.01 29.75
PatchNet [59] 78.15 41.50
IADG [78] 78.55 26.27
@@
LLaVA [38] 16.39
Qwen-VL [4] 16.55
Minigpt4 [79] 19.51
Bunny [20] 27.03
Bunny (Fine-tuned) [20] 94.43
@@
LLaVA [38] 45.05 31.75 23.42 17.80 30.51 25.52 37.84 50.11
Minigpt4 [79] 17.85 7.85 3.53 1.94 27.54 21.88 33.86 50.00
Qwen-VL [4] 20.92 14.53 11.01 8.77 21.45 12.49 47.64 41.49
Bunny [20] 33.64 27.12 22.65 19.33 36.74 19.12 50.68 39.73
Bunny(fine-tuned) [20] 89.57 86.96 84.91 81.29 80.15 51.64 98.56 1.16

### v2

@@
Abstract
Face anti-spoofing (FAS) is crucial for protecting facial recognition systems from presentation attacks. Previous methods approached this task as a classification problem, lacking interpretability and reasoning behind the predicted results. Recently, multimodal large language models (MLLMs) have shown strong capabilities in perception, reasoning, and decision-making in visual tasks. However, there is currently no universal and comprehensive MLLM and dataset specifically designed for FAS task. To address this gap, we propose FaceShield, a MLLM for FAS, along with the corresponding pre-training and supervised fine-tuning (SFT) datasets, FaceShield-pre10K and FaceShield-sft45K. FaceShield is capable of determining the authenticity of faces, identifying types of spoofing attacks, providing reasoning for its judgments, and detecting attack areas. Specifically, we employ spoof-aware vision perception (SAVP) that incorporates both the original image and auxiliary information based on prior knowledge. We then use an prompt-guided vision token masking (PVTM) strategy to random mask vision tokens, thereby improving the model’s generalization ability. We conducted extensive experiments on three benchmark datasets, demonstrating that FaceShield significantly outperforms previous deep learning models and general MLLMs on four FAS tasks, i.e., coarse-grained classification, fine-grained classification, reasoning, and attack localization. Our instruction datasets, protocols, and codes will be released at https://github.com/Why0912/FaceShield.
Figure 1: FaceShield Multi-task Response Demonstration. This figure shows the model’s performance on four tasks: coarse-grained classification (real vs. spoofed faces), fine-grained classification (specific attack types like print attacks), reasoning (explaining spoofing using features such as lighting and symmetry), and localization (detecting spoofed regions). It highlights FaceShield’s ability to handle diverse, complex questions accurately.
Introduction
@@
Existing deep learning FAS models can be categorized into two types: vision-based methods and vision-language-based methods. As shown in Fig. 2(a), vision-based methods rely solely on image data (e.g., RGB, Depth, Infrared(IR)) and binary labels to train CNNs (yu2020searching; yu2021dual; wang2025pnss) or ViTs (9484333; yu2023rethinkingvisiontransformermasked; cai2025rehearsal) for FAS. While they can achieve satisfactory results against known attack types and environments, these methods are prone to overfitting on spurious correlations and lack strong extrapolation capabilities. As illustrated in Fig. 2(b), Vision-language-based methods do not use binary labels but instead train CLIPs with image-text pairs (Srivatsan_2023_ICCV; liu2024fmclip; mu2024tegdgtextuallyguideddomain; lin2025reliable). The text labels in these methods provide more domain-agnostic information, enhancing models’ generalization capability. Although these FAS models demonstrate some recognition capabilities, they still face challenges such as limited generalization ability, poor interpretability, and a lack of capability for fine-grained localization of attack regions.
Recently, MLLMs have shown remarkable capabilities across various visual tasks (ye2025cat), such as remote sensing (zhang2024earthgpt; kuckreja2023geochat; 2402.02544), medical imaging (li2023llavamed; Sun2024STLLaVAMedSL), and deepfake detection  (xu2024fakeshieldexplainableimageforgery; huang2024ffaa). By leveraging the general capabilities of language foundation models alongside the visual information extracted by vision towers, these specialized MLLMs integrate perception, reasoning, and decision-making within a single model. Regarding the FAS task, SHIELD (shi2025shield) conducted extensive evaluations on existing general-purpose MLLMs, revealing that their performance on FAS tasks still has room for improvement. (zhang2025interpretable) introduced a model capable of performing classification and description attack type. However, the model is limited in its ability to handle more nuanced tasks, such as identifying specific attack types, reasoning, and localizing spoofed areas. These limitations underscore the broader challenges in training FAS MLLMs, including: (1) a lack of pretraining and SFT datasets specific to FAS tasks, (2) the need to extend traditional FAS tasks to fully exploit MLLM capabilities, and (3) the difficulty for general-purpose vision towers to capture the subtle distinctions between real faces and PAs, unlike with natural images.
@@
Figure 3: Construction pipeline of our proposed instruction datasets (i.e., FaceShield-pre10K and FaceShield-sft45K).The initial datasets (WMCA, PADISI, SiW-Mv2) are combined to form a uni-class dataset covering 12 spoofing types, with selected images annotated for visual grounding. Using MLLM with structured prompts, we generate two datasets: a pretraining dataset and an SFT dataset divided into four tasks (coarse-grained classification, fine-grained classification, reasoning, and localization). The pretraining data is filtered by CLIP for similarity, producing the FaceShield-pre10k dataset. SFT data undergoes multi-level filtering (LLM-based, keywords, and human reviews), followed by augmentation, resulting in the FaceShield-sft45k dataset. Additional details can be found in the appendix.
Related Work
Face Anti-Spoofing. Early FAS methods primarily used CNNs (yu2020searching) and ViTs (9484333), incorporating auxiliary cues such as reflection (icassp22multiple), depth (cvpr18aux), and rPPG (yu2019remote) for live/spoof classification. While some fuse multi-modal inputs (e.g., RGB, IR, depth)(yu2024rethinking), they often suffer performance drops in unseen domains due to domain shifts(tpami23survey). To enhance generalization, domain-generalized FAS methods (tpami23survey) employ techniques such as adversarial training (ijcv23adversarial), feature disentanglement (cvpr22ssan), one-class learning (cvpr24oneclass), meta-learning (qin2021meta), data augmentation (ijcv24augmentation), and data synthesis (aaai22generation) to model domain shifts and separate domain-invariant content from domain-specific styles. Additionally, domain or gradient alignment (cvpr24gradient) aids in learning generalized decision boundaries for more robust and domain-invariant spoof detection.
Recently, vision-language models (VLMs), especially CLIP (clip), have been explored for FAS (flip; liu2024fmclip). These models use textual descriptions of live and spoof faces to guide classification, offering stronger generalization by leveraging the powerful visual representations learned during pre-training (liu2024fmclip).
However, most FAS methods lack interpretability, particularly in providing language-based explanations. A recent work (zhang2025interpretable) collected 12 datasets and used instruction tuning to train an MLLM for attack classification, but it struggles to integrate classification, localization, and reasoning tasks for real-world deployment. Additionally, existing FAS datasets (eccv25tffas; cvpr24cfpl) lack reasoning annotations, hindering the development of more comprehensive and interpretable models.
Multimodal Large Language Models. MLLMs such as GPT-4V (gpt4), LLaVA (liu2023llava), and Bunny (he2024bunny) have shown strong performance across general domains (li2023seedbench2benchmarkingmultimodallarge; huang2024aesbenchexpertbenchmarkmultimodal). Task-specific MLLMs have also been developed for various vision tasks, including remote sensing (zhang2024earthgpt; kuckreja2023geochat), medical imaging (li2023llavamed; Sun2024STLLaVAMedSL), aesthetic assessment (huang2024aesexpert), deepfake detection (xu2024fakeshieldexplainableimageforgery; huang2024ffaa), and visual grounding (wei2023lenna; peng2023kosmos). Unlike traditional models, MLLMs unify multiple tasks within a single framework and generalize well to unseen domains. For example, FakeShield (xu2024fakeshieldexplainableimageforgery) introduces an explainable multi-task model for image forgery detection and localization by leveraging visual-textual clues at both global and pixel levels.
However, current MLLMs lack specific knowledge about face attacks, leading to limited performance on FAS tasks. To address this, we introduce dedicated FAS datasets that enrich MLLMs with domain-specific knowledge. The proposed SAVP module enhances spoof discrimination by guiding the model with semantic attack priors, while PVTM improves feature generalization across varied domains.
FaceShield-pre10K and FaceShield-sft45K
Dataset Collection
Figure 4: Proposed model architectures. (a) Proposed model with Spoof-Aware Vision Perception (SAVP). (b) Proposed model with SAVP and Prompt-Guided Vision Token Masking (PVTM). (c) Details about PVTM.
@@
Fig. 3(a) illustrates the annotation process for existing FAS datasets. Based on the class types from WMCA (W) (8714076), PADSIS (P) (Rostami_2021_ICCV), and SiW-Mv2 (S) (xiaoguo2022MDFAS), we unify the annotation categories into 12 types: Bonafide, Fakehead, Print, Glasses, Replay, Paper mask, Flexible mask, Rigid mask, Partial Eye, Partial Mouth, Makeup, and Tattoo. We re-annotate all images at both image- and region-levels, resulting in 12,091 images with class labels and 3,139 images with bounding box annotations. The detail information is shown in Table 1.
Instruction Construction
As shown in Fig. 3(b), we construct two instruction datasets using Bunny-Llama-3-8B-V (he2024bunny). A system prompt containing class labels is used to guide the MLLM assistant in generating question-answer (QA) pairs, based on the task type and few-shot examples.
@@
For the pretraining dataset FaceShield-pre10K, we generate image descriptions only, without task instructions. Pairs with a CLIP (radford2021learning) similarity score below 15% are filtered out to ensure quality.
Algorithm 1 Construction of FaceShield-sft45K Dataset
1:Image dataset $\mathcal{D}_{img}$ from WMCA, SiW-Mv2, PADISI;
2:Multimodal LLM (MLLM); Task set $\mathcal{T}$.
3:Processed QA dataset $\mathcal{D}_{QA}$.
4:Initialize $\mathcal{D}_{QA}\leftarrow\emptyset$
5:for image $I\in\mathcal{D}_{img}$ do
6:  for task $t\in\mathcal{T}$ do
7:   Generate task-specific prompt $\mathcal{P}_{t}$ using Ground Truth (GT), requirements and examples
8:   Generate QA pair $(q,a)$ using MLLM: $(q,a)=\text{MLLM}(\mathcal{P}_{t},I)$
9:   if QA pair $(q,a)$ satisfies validation criteria then
10:     Add $(q,a)$ to $\mathcal{D}_{QA}$
11:   end if
12:  end for
13:end for
14:Filter low-quality QA pairs in $\mathcal{D}_{QA}$:
15:  Remove pairs with low semantic consistency or invalid content
16:  Perform human review to refine remaining pairs
17:Augment $\mathcal{D}_{QA}$ by generating rephrased QA pairs:
18:for QA pair $(q,a)\in\mathcal{D}_{QA}$ do
19:  Generate additional pairs $(q^{\prime},a^{\prime})$ using rephrasing strategies
20:  Add $(q^{\prime},a^{\prime})$ to $\mathcal{D}_{QA}$
21:end for
22:return $\mathcal{D}_{QA}$
For the instruction-tuning dataset FaceShield-sft45K, as shown in Alg. 1, MLLM-generated QA pairs are filtered using both manual and keyword-based strategies. The resulting high-quality seed set is then diversified using LLaMA3 (dubey2024llama) to enhance linguistic richness and dialogue ability.
Table 1: Datasets and QA statistics
Datasets Attack Types Annotations QA_Count
FaceShield-pre10K - Image-text pairs 9297
FaceShield-sft45K Unified-attack(11 types) QA + bbox 45662
The dataset covers four tasks: (1) Coarse-grained classification, which predicts whether a face is real or spoofed; (2) Fine-grained classification, which identifies specific PA types beyond binary classification; (3) Reasoning, which provides explanations and justifications before making a judgment; (4) Attack localization, which outputs coordinates of attack regions if spoofing is detected.
FaceShield
@@
Our goal is to train a FAS task-specific MLLM with two main objectives: 1) Enhance the visual encoder’s ability to extract features from real faces and presentation attacks, and 2) Utilize the extensive knowledge stored in the LLM to improve the model’s generalization capabilities when facing unknown domains.A naive training approach involves direct pre-training and SFT using RGB images and constructed QA data. However, the high similarity between real faces and PAs in RGB appearance poses significant challenges to this method. As shown in Fig. 4(a), Spoof-Aware Vision Perception (SAVP) combines images preprocessed based on prior knowledge, by extracting predefined local descriptor operators (yu2024rethinking), with the original RGB image. Our complete model framework is shown in Fig. 4(b), RGB images and the extracted local descriptor images are fed into the vision encoder to extract vision token $V_{RGB}$ and $V_{SAV}$, respectively. These features are then processed through the Prompt-Guided Vision Token Masking (PVTM) module, which extracts highly generalizable vision tokens. These tokens are sent to a projector to align with text prompt token $P$ to produce $V_{align}$, which is then fed into the language foundation model for inference result $\mathcal{Y}$ as follows:
@@
Spoof-Aware Vision Perception
Bonafide faces and PAs lack distinct discriminative features in RGB-based appearance space, whereas local descriptors (yu2020searching; yu2024rethinking; xie2024fusionmamba) extracted through image preprocessing can enhance their subtle live/spoof clues. As shown in Fig. 4(a), we extract features from the original images using Local Binary Pattern (LBP) (pietikainen2010local), Gray, and Histogram of Oriented Gradients (HOG) (dalal2005histograms), and concatenate them. LBP and Gray-specific computations are as follows:
@@
[…]
@@
Prompt-Guided Vision Token Masking
To further enhance the alignment between visual features and text prompts, and alleviate overfitting on spurious correlations, we leverage text prompts to guide token selection after the visual encoder. As shown in Fig. 4(c), the text tokens extracted from the prompt are pooled (song2024moresimpleeffectivetoken), and then their similarity with all visual tokens is calculated. We assume that visual tokens with higher similarity are more relevant to subsequent tasks and even for the same image, the important visual tokens may not be consistent across different tasks. The calculation of similarity between visual tokens $V_{i}$ and text prompt tokens $P$ is as follows:
@@
Training Details
We adopt a two-stage training strategy. In the first stage (pretraining), we perform continual pretraining to align visual embeddings from a pretrained vision encoder with text embeddings using FaceShield-pre10K. In the second stage (supervised fine-tuning, SFT), we apply visual instruction tuning on FaceShield-sft45K to fully exploit the MLLM’s capabilities across domain-specific multimodal tasks.
For LLM adaptation, we apply LoRA (hu2021loralowrankadaptationlarge) with a rank of 128 and a scaling factor of 256 for each transformer block. Cross-entropy loss is used for next-token prediction in both stages. During pretraining, we update only the vision projector and the PVTM module for one epoch. In the SFT stage, we fine-tune the LoRA layers in the LLM along with the vision projector.
Experiments and Results
Protocols and Evaluation Metrics
We use 10% of each source dataset in FaceShield-sft45K to construct three test subsets: W, S, and P. For coarse-grained classification we perform both intra- and cross-dataset evaluations. For fine-grained classification, reasoning and attack localization, we conduct intra-dataset evaluation only. In intra-dataset testing, models are trained on all source data and evaluated on the combined test sets (W&S&P). For cross-dataset testing, two datasets are used for training (including pretraining and SFT), and the remaining one for testing (e.g., training on W and S, testing on P).For coarse-grained classification, fine-grained classification, and reasoning, we report Half Total Error Rate (HTER) (yu2022deep) and Accuracy (ACC). For reasoning, we further evaluate reasoning quality using BLEU (papineni2002bleu), ROUGE-L (lin2004rouge), and METEOR (banerjee2005meteor). For attack localization, we report AP@40 and AP@50.
Implementation Details
We use Siglip (zhai2023sigmoid) as the visual encoder and Phi-3 (abdin2024phi) as the language foundation model. PVTM retains the top 10% of the most important tokens and randomly masks 5% of the tokens in the remaining 90%. Adam optimizer is used in the pretrain stage with a learning rate of $5\times 10^{-4}$. As for SFT stage, we decrease the learning rate to $2\times 10^{-4}$. All experiments are conducted on a single NVIDIA A100 GPU. Each experiment is repeated 10 times on the model, and the final results are reported as the mean ± standard deviation.
Comparison with Existing Methods
Coarse-Grained Classification Task
For the coarse-grained classification task, we compare FaceShield with state-of-the-art FAS methods  (he2016deep; Wang_2022_CVPR; zhou2022learning; zhou2023instance) and open-source MLLMs (liu2023llava; Qwen-VL; zhu2023minigpt; he2024bunny).
@@
ResNet (he2016deep) 97.55 2.32
PatchNet (Wang_2022_CVPR) 98.22 1.78
CoOp (zhou2022learning) 98.73 1.27
@@
LLaVA (liu2023llava) 65.54 27.76
Qwen-VL (Qwen-VL) 51.94 38.70
Minigpt4 (zhu2023minigpt) 26.86 65.50
Bunny (he2024bunny) 81.20 17.87
Bunny (fine-tuned) (he2024bunny) 98.23 1.52
@@
W & S → P
ResNet (he2016deep) 46.12 50.00
PatchNet (Wang_2022_CVPR) 77.18 22.87
IADG (zhou2023instance) 72.96 27.01
FAS-AUG (ijcv24augmentation) 91.7 7.3
@@
W & P → S
ResNet (he2016deep) 53.36 49.16
PatchNet (Wang_2022_CVPR) 56.16 45.37
IADG (zhou2023instance) 57.20 42.81
FAS-AUG (ijcv24augmentation) 88.2 11.7
@@
S & P → W
ResNet (he2016deep) 74.01 29.75
PatchNet (Wang_2022_CVPR) 78.15 41.50
IADG (zhou2023instance) 78.55 26.27
FAS-AUG (ijcv24augmentation) 87.9 13.1
@@
Table 4: Cross-dataset result on CASIA-MFSD and Replay-Attack.
Test Dataset ACC(%) $\uparrow$ HTER(%) $\downarrow$
S & P & W $\rightarrow$ CASIA-MFSD 90.59 6.37
S & P & W $\rightarrow$ Replay-Attack 82.42 20.07
@@
LLaVA (liu2023llava) 16.39
Qwen-VL (Qwen-VL) 16.55
Minigpt4 (zhu2023minigpt) 19.51
Bunny (he2024bunny) 27.03
Bunny (Fine-tuned) (he2024bunny) 94.43
@@
LLaVA (liu2023llava) 45.05 31.75 23.42 17.80 30.51 25.52 37.84 50.11
Minigpt4 (zhu2023minigpt) 17.85 7.85 3.53 1.94 27.54 21.88 33.86 50.00
Qwen-VL (Qwen-VL) 20.92 14.53 11.01 8.77 21.45 12.49 47.64 41.49
Bunny (he2024bunny) 33.64 27.12 22.65 19.33 36.74 19.12 50.68 39.73
Bunny(fine-tuned) (he2024bunny) 89.57 86.96 84.91 81.29 80.15 51.64 98.56 1.16